OpenStudy (anonymous):

Multiply both sides of r=-6sinø by r and use the fact that r^2=x^2+y^2 to rewrite this equation in terms of x and y

OpenStudy (anonymous):

Hint: \[y=rsin \theta\]

Directrix (directrix):

r=-6sinø r * r = r * (-6sinø) r^2 = -6 * r sinø Use the hint above. r^2 = -6 * y Substitute for r^2 as given in the instructions. And then ... @DurableToaster

Directrix (directrix):

This mess --> ������������������) is the angle phi.

Directrix (directrix):

@DurableToaster